A sphere has radius R 16.1 measured in cm. What is the volume of this sphere measured in in3? Hint: The volume of a sphere is given by VR
Answer #1

Volume, V = (4/3) pi x R^3

R = 16.1 cm = 6.339 in

V = (4/3) pi x (6.339)^3 = 1066.97 in^3

Comment in case any doubt please rate my answer ....